Grade 9Mechanics

A gun of mass M fires a bullet of mass m with a Kinetic energy E . The velocity of the recoil of the gun is

Somesh
7 Years agoGrade 9

2 Answers

Arun
7 Years ago
Momentum before firing = 0 
Momentum after firing = mv + MV

According to law of conservation of momentum, 

mv + MV = 0
Recoil velocity V = - mv /M
Khimraj
7 Years ago
mass of bullet = m
velocity of bullet v = \sqrt{2E/m}
mass of gun = M
recoil velocity of gun V=?
apply the law of conservation of momentum
mv + MV = 0
V = -mv/M = -(\sqrt{2Em})/M
Hope it clears............................